OpenStudy (anonymous):

What are the lengths of the missing side of a 30° -60° -90° triangle if the longer leg of the triangle is 18 centimeters?

OpenStudy (aaronq):

 \(sin\theta=\dfrac{opposite}{hypoteneuse}\) \(cos\theta =\dfrac{adjacent}{hypotenuese} \) \(tan \theta=\dfrac{opposite}{adjacent}\)